OpenStudy (anonymous):

PLEASE HELP, I need the answer! it would make a huge difference. 6 – 3/4x + 1/3 = 1/2x + 5

OpenStudy (rhr12):

Do u need the value of x?

OpenStudy (anonymous):

Yes here are the option: a.2 b.3 c.6 d.12

OpenStudy (anonymous):

*options

OpenStudy (anonymous):

@rhr12 it'd e great if you could help

OpenStudy (rhr12):

Okay. Hold on

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

\[6-\frac{ 3 }{ 4 }x+\frac{ 1 }{ 3 }=\frac{ 1 }{ 2 }x+5\]\[6\frac{ 1 }{ 3 }-\frac{ 3 }{ 4 }x=\frac{ 1 }{ 2 }x+5\]\[1\frac{ 1 }{ 3 }-\frac{ 3 }{ 4 }x=\frac{ 1 }{ 2 }x\]\[2\frac{ 2 }{ 3 }-\frac{ 6 }{ 4 }x=x\]\[2\frac{ 2 }{ 3 }=\frac{ 10 }{ 4 }x\]\[\frac{ 8 }{ 3 } = \frac{ 10 }{ 4 }x\]\[8=\frac{ 30 }{ 4 }x\]\[32=30x\]

OpenStudy (anonymous):

but this answer ain't one of the options

OpenStudy (rhr12):

I got 16/3

OpenStudy (anonymous):

Thanks for the help! but yes it's not there... I just re-read it. The question wants to know which number can each term of the equation be multiplied by to eliminate the fractions before actually solving it.

OpenStudy (anonymous):

wait i made a typo in the working...

OpenStudy (anonymous):

Sorry for the mistake, do you know the answer?

OpenStudy (rhr12):

Is it (6-3)/x +1/3 = 1/2x+5?

OpenStudy (rhr12):

Or 6-(3/x)?

OpenStudy (anonymous):

The answer? Its one of these: a.2 b.3 c.6 d.12

OpenStudy (rhr12):

Which one?

OpenStudy (anonymous):

No idea, I should know but.... I don't :(

OpenStudy (anonymous):

plug in every option and try :)

OpenStudy (anonymous):

Ya, I guess I will...Thanks guys! :)

OpenStudy (rhr12):

The problem is braces. U didn't indicate where it is. That is so confusing

OpenStudy (anonymous):

Let me send you the whole question

OpenStudy (anonymous):

Which number can each term of the equation be multiplied by to eliminate the fractions before solving? 6 – 3/4 x + 1/3 = 1/2 x + 5 a.2 b.3 c.6 d.12

OpenStudy (anonymous):

d

OpenStudy (anonymous):

its asking what of this number you can multiply on both sides and not get any fractions

OpenStudy (rhr12):

D is not correct i guess

OpenStudy (anonymous):

@saseal It's fine, but thanks! @rhr12 Thank you for trying :)

OpenStudy (rhr12):

Do u get the correct answer?

OpenStudy (rhr12):

I think it's 3

OpenStudy (anonymous):

Hopefully, I'll let you know. Sucks because they don't let you know (the program I am in) it just goes toward your grade.

OpenStudy (anonymous):

i wish my school does this kind of test

OpenStudy (rhr12):

Aight don't choose three. Take your own guess and best of luck. I took online geo final yesterday and i took guess for more than 30 questions but i get 81. But i don't know how. There were 51 questions. Believe yourself and take ur guess. 😝

OpenStudy (anonymous):

@saseal great online class but long videos, guess I should put more attention :( @rhr12 Thanks a lot I really appreciate the help, ya guessung comes in handy

OpenStudy (anonymous):

*guessing

OpenStudy (anonymous):

i usually take online classes from edx but they are at calculus and above level. the reason why i tell you 12 is theres 3 fractions in this equation , 3/4, 1/3, 1/2. only 12 is the LCM of this 3 denominator

OpenStudy (anonymous):

Calculus, that's great! I bet you like Math

OpenStudy (anonymous):

nah, but its a requirement in my civil engineering course

OpenStudy (anonymous):

College? University?

OpenStudy (anonymous):

a Pre-Uni tech school

OpenStudy (anonymous):

we can skip the 1st year of civil engineering in a local uni in my country and continue to degree with their diploma

OpenStudy (anonymous):

Awesome, sounds great. Less is better, I am assuming it will get you there sooner?

OpenStudy (anonymous):

Obviously... :)

OpenStudy (anonymous):

its 1 year longer than college to uni but less stressful and you get to actually use what you learned

OpenStudy (anonymous):

its more fun

OpenStudy (anonymous):

our college is 2 years and tech school is 3 years

OpenStudy (rhr12):

@saseal what state u r in?

OpenStudy (anonymous):

Not that bad, 1 more year makes a great difference. You could spend that year doing other things but instead using it there is best. And like you said things you will use and it's fun... it's worth your time :)

OpenStudy (anonymous):

singapore, asia

OpenStudy (anonymous):

Very cool!

OpenStudy (rhr12):

I'm kinda asian. I mean my parents are from south asia

OpenStudy (rhr12):

@saseal

OpenStudy (anonymous):

thats cool

OpenStudy (rhr12):

Okay. Bye. I better give u a medal

OpenStudy (anonymous):

bai too

OpenStudy (anonymous):

Bye, got to go. Thanks again!

OpenStudy (anonymous):

yw